7. Calculate the fraction of the starting quantity of A that will be used up after 60 s. Given the reaction below which is found to be the first order in A and t1/2=40s A→B+C
Answer #1

Given:

Half life = 40 s

use relation between rate constant and half life of 1st order reaction

k = (ln 2) / k

= 0.693/(half life)

= 0.693/(40)

= 1.733*10^-2 s-1

we have:

[A]o = 100 [Let initial concentration be 100]

t = 60.0 s

k = 1.733*10^-2 s-1

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln[A] = ln(1*10^2) - 1.733*10^-2*60

ln[A] = 4.605 - 1.733*10^-2*60

ln[A] = 3.566

[A] = e^(3.566)

[A] = 35.4

35.4 of 100 is remaining.

Amount used up is 100 - 35.4 = 64.6 which is 64.6 %

Answer: 64.6 %